OpenStudy (anonymous):

Find the radius of a circle whose circumference is 31.4 centimeters.

zepdrix (zepdrix):

Hey there! :) So we have a formula that relates Circumference to the Radius of a circle.

OpenStudy (anonymous):

?

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{C}=2\pi r\]Let's rewrite our pi as 3.14 so it's easier to work with:\[\Large\rm \color{orangered}{C}=2\cdot 3.14 r\]Multiplying the 2 and 3.14 together gives us:\[\Large\rm \color{orangered}{C}=6.28 r\]And they told us that the circumference of this particular circle is \(\Large\rm \color{orangered}{31.4}\).

zepdrix (zepdrix):

So our equation is:\[\Large\rm \color{orangered}{31.4}=6.28 r\]From here we need to solve for r, understand how to do that?

OpenStudy (anonymous):

Find the area of a circle whose circumference is 31.4 centimeters. So you do 3.14 times 31.4

zepdrix (zepdrix):

No. We've simplified our equation to 31.4 = 6.28r See how 6.28 is `multiplying` the r? To isolate the r, we do the inverse of `multiplication`. So we'll divide each side by 6.28.

zepdrix (zepdrix):

So then you turn to your trusty calculator to finish things up for you. r = 31.4 / 6.28

OpenStudy (anonymous):

1.Find the radius of a circle whose circumference is 31.4 centimeters. =6.28 Find the area of a circle whose circumference is 31.4 centimeters =5

OpenStudy (anonymous):

2. Find the area of a circle whose circumference is 31.4 centimeters =5

zepdrix (zepdrix):

Hmmm no :( If you followed my instructions, you should get that the radial value is `r=5`. The question after that uses this information to calculate area. So our r=5 Our formula for area is A=3.14 r^2 So plug in the r value that found in the first question. A=3.14 (5)^2 and simplify.

OpenStudy (anonymous):

I'm very confused so number 1 is. 5

zepdrix (zepdrix):

yes

OpenStudy (anonymous):

Then what would number 2 be